# SFTMix: Elevating Language Model Instruction Tuning with Mixup Recipe

### v2

@@ 4.2 Domain-Specific SFT
In domain-specific SFT for healthcare, we train Llama and Mistral on the MedAlpaca-263K medical conversation dataset (Han et al., 2023) using either NTP or SFTMix for two epochs, while keeping the remaining hyperparameters as specified in §4.1. We then assess their performance on four healthcare-related question-answering benchmarks: MedQA (Jin et al., 2021), its five-option variant MedQA-5, PubMedQA (Jin et al., 2019), and MedMCQA (Pal et al., 2022). We adopt the three-shot evaluation setting from Labrak et al. (2024) and report the results over three evaluation rounds. Additionally, we include prior biomedical LLMs of similar sizes, including MedAlpaca-7B (Han et al., 2023), PMC-LLaMA-7B (Wu et al., 2024), BioMedGPT-LM-7B (Luo et al., 2023), Meditron-7B (Chen et al., 2023), and BioMistral-7B (Labrak et al., 2024), as reference models for comparison.
@@
5 Analysis
Inspired by Burns et al. (2024), we investigate the generalization of training dynamics from a weaker reference LLM to a stronger instruction-tuning LLM. Specifically, we identify training dynamics with a weaker reference LLM, Gemma-2B (Team et al., 2024), to divide Alpaca-52K into a confident subset and an unconfident subset of equal size. These subsets are then fed into the Mixup regularization when instruction-tuning Llama.
In Table 3, this alternative approach yields comparable scores on MT-Bench and AlpacaEval-2 to the default SFTMix recipe, which uses the same LLM for both training dynamics and Mixup-based instruction tuning. This finding aligns with the weak-to-strong generalization reported by Burns et al. (2024) and highlights the potential for scaling SFTMix to even stronger LLMs.
@@ 5.2 Training Dynamics-Based Confidence Is Crucial for Performing Mixup
We now explore whether we can substitute training dynamics-based confidence with known data quality. To test this hypothesis, we replace half of the original responses in Alpaca-52K with higher-quality GPT-4-generated versions (Peng et al., 2023), forming the “High” subset, while referring to the remaining lower-quality original responses as “Low”. We then train Llama using three approaches: (1) NTP on High, (2) NTP on the combined High $+$ Low dataset, and (3) NTP on High $+$ Low with the Mixup regularization applied between them.
@@
5.3 Incorporating Mixup as a Regularization Is More Effective
@@ 5.4 SFTMix Effectively Utilizes Entire Instruction-Tuning Datasets
5.5 SFTMix Integrates Well with Data Selection Methods
Although SFTMix performs effectively with the uncurated Alpaca-52K in §4.1, it can be seamlessly integrated with various data selection methods. Here, we first select $1{,}000$ high-quality examples from Alpaca-52K using either AlpaGasus (Chen et al., 2024), which grades responses with proprietary LLMs, or Long (Zhao et al., 2024), which chooses the longest responses. We then apply either NTP or SFTMix to the selected examples.
As shown in Table 7, instruction-tuning on the subset selected by AlpaGasus achieves performance similar to training on the full uncurated dataset, while using the longest examples leads to slightly better results than both alternatives. Nevertheless, applying SFTMix alongside both data selection methods still yields substantial improvements over the conventional NTP baseline. This suggests that integrating SFTMix with existing data selection strategies (Albalak et al., 2024; Wang et al., 2024) could further enhance performance in LLM instruction tuning.
@@ 6 Conclusion
In this paper, we propose SFTMix, a novel recipe for elevating LLM instruction tuning, without relying on well-curated SFT datasets. We observe that LLMs exhibit uneven confidence distributions across the semantic representation space. Because confident examples are prone to overfitting and unconfident ones are harder to generalize, we argue that data with different confidence levels should play distinct roles in instruction tuning. Building on this motivation, we first derive an LLM’s confidence from its training dynamics and partition the SFT dataset into confident and unconfident subsets of equal size. We then interpolate between these subsets to bridge the confidence gap and introduce a Mixup-based regularization to facilitate learning on these additional, interpolated examples. In this way, SFTMix propagates supervision signals across confidence regions and encourages linearity over a smoother decision boundary between these regions. This process attenuates the risk of overfitting in confident examples and enhances generalization in unconfident ones. Extensive empirical results in both instruction-following and domain-specific SFT tasks demonstrate that SFTMix outperforms the conventional NTP paradigm across various LLM families and SFT datasets of different sizes and qualities. Our in-depth analyses across six directions further highlight that SFTMix integrates seamlessly with data selection methods, adapts well to compute-constrained scenarios, and scales effectively to broader applications.
7 Limitation
Due to computational constraints, we do not apply SFTMix to LLM pre-training or instruction-tune larger LLMs using this recipe. Additionally, while we validate SFTMix on instruction-following and domain-specific SFT tasks, our evaluations are limited to English-based benchmarks, and we do not investigate its impact on the safety alignment (Liang et al., 2023; Liu et al., 2023) of LLMs. We leave these explorations for future work.
Appendix A How Does SFTMix Affect Gradient Descent?
In SFTMix, we linearly interpolate examples with different confidence levels and introduce a Mixup-based regularization to support learning on these additional, interpolated examples. Here, we analyze how this regularization influences the direction of gradient descent in the original NTP paradigm. For notational simplicity, we focus on the cross-entropy $H(\tilde{\mathbf{Y}},\sigma(\tilde{\mathbf{Z}}^{\top}\mathbf{W}))$ between the interpolated one-hot encoding vector $\tilde{\mathbf{Y}}$ and the corresponding interpolated representation $\tilde{\mathbf{Z}}$ from the last transformer layer of the target instruction-tuning LLM.
Let $\tilde{\mathbf{S}}=\tilde{\mathbf{Z}}^{\top}\mathbf{W}$, then the gradient of $H$ w.r.t. $\tilde{\mathbf{S}}$ is simply
Using the chain rule, we have
$\nabla_{\mathbf{W}}H(\tilde{\mathbf{Y}},\sigma(\tilde{\mathbf{Z}}^{\top}\mathbf{W}))=\tilde{\mathbf{Z}}^{\top}\left(\sigma(\tilde{\mathbf{Z}}^{\top}\mathbf{W})-\tilde{\mathbf{Y}}\right).$
In other words, the gradient from the regularization with the interpolated example does not decompose into a weighted sum of the gradients from the NTP loss applied to the corresponding confident and unconfident examples. As a result, the Mixup-based regularization modifies, rather than simply reweights, the gradient descent direction in NTP by incorporating these interpolated examples.
Appendix B Additional Experiment Results
B.1 Hyperparameter Search for $\mu$
@@
In SFTMix, we use the hyperparameter $\mu$ to control the regularization effect in the training loss $\ell_{\text{SFTMix}}$. To explore its impact, we experiment with $\mu\in\{0.1,0.2,0.5\}$ by instruction-tuning Llama on Alpaca-52K. As shown in Table 9, $\mu=0.2$ achieves the highest performance across most metrics, except for the multi-turn conversational ability in MT-Bench. Therefore, we set $\mu=0.2$ as the default for the experiments in §4 and §5.
@@
Table 9: Hyperparameter search for $\mu\in\{0.1,0.2,0.5\}$. We set $\mu=\underline{0.2}$ as the default in SFTMix, as it achieves the highest performance across most metrics.
B.2 Human Evaluation on AlpacaEval-2

### v3

@@ 4.2 Domain-Specific SFT
In healthcare-specific SFT, we train Llama and Mistral on the MedAlpaca-263K medical dataset (Han et al., 2023) using either NTP or SFTMix for two epochs, keeping other hyperparameters as in §4.1. We assess their performance on four healthcare-related question-answering benchmarks: MedQA (Jin et al., 2021), its five-option variant MedQA-5, PubMedQA (Jin et al., 2019), and MedMCQA (Pal et al., 2022). Following Labrak et al. (2024), we adopt a three-shot setting and report the mean accuracy over three evaluation runs. For comparison, we also include prior biomedical LLMs of similar size: MedAlpaca-7B (Han et al., 2023), PMC-LLaMA-7B (Wu et al., 2024), BioMedGPT-LM-7B (Luo et al., 2023), Meditron-7B (Chen et al., 2023a), and BioMistral-7B (Labrak et al., 2024).
@@
5 Analysis
Inspired by Burns et al. (2024), we explore whether training dynamics from a weaker reference LLM transfer to a stronger one. Specifically, we use Gemma-2B (Team et al., 2024) to split Alpaca-52K into confident and unconfident subsets, then use them for Mixup regularization when tuning Llama.
In Table 3, this alternative approach yields comparable scores on MT-Bench and AlpacaEval-2 to the default SFTMix recipe, which uses the same LLM for both training dynamics and Mixup-based instruction tuning. This finding aligns with the weak-to-strong generalization Burns et al. (2024) and suggests that confidence scores precomputed from weaker LLMs can be reused for stronger models, amortizing the computational cost over time.
@@ 5.2 Training Dynamics-Based Confidence Is Crucial for Performing Mixup
We now explore whether we can substitute training dynamics-based confidence with known data quality. To test this hypothesis, we replace half of the original responses in Alpaca-52K with higher-quality GPT-4-generated versions (Peng et al., 2023), forming the “High” subset, while referring to the remaining lower-quality original responses as “Low”. We then train Llama using three approaches: (1) NTP on High, (2) NTP on the combined High ++ Low dataset, and (3) NTP on High ++ Low with the Mixup regularization applied between them.
@@
5.3 Incorporating Mixup as a Regularization Is More Effective
@@ 5.4 SFTMix Effectively Utilizes Entire Instruction-Tuning Datasets
5.5 SFTMix Integrates Well with Data Selection Methods
Although SFTMix performs effectively with the uncurated Alpaca-52K in §4.1, it can be seamlessly integrated with various data selection methods. Here, we first select 1,0001{,}000 high-quality examples from Alpaca-52K using either AlpaGasus (Chen et al., 2024a), which grades responses with proprietary LLMs, or Long (Zhao et al., 2024), which chooses the longest responses. We then apply either NTP or SFTMix to the selected examples.
As shown in Table 7, applying SFTMix to the uncurated dataset outperforms NTP with either data selection strategy. Instruction tuning on the AlpaGasus-selected subset matches performance on the full dataset, while using the longest examples performs slightly better. Nevertheless, combining SFTMix with either method yields substantial improvements over the NTP baseline. This suggests that integrating SFTMix with existing data selection strategies (Albalak et al., 2024; Wang et al., 2024) could further enhance performance in LLM instruction tuning.
@@ 6 Conclusion
In this paper, we propose SFTMix, a novel recipe for elevating LLM instruction tuning. We observe that LLMs exhibit uneven confidence distributions across the semantic space, and argue that data with different confidence levels should play distinct roles in tuning. To this end, we partition an SFT dataset into confident and unconfident subsets, interpolate them to bridge the confidence gap, and introduce a Mixup-based regularization to facilitate learning. Extensive experiments show that SFTMix outperforms the conventional NTP paradigm across diverse LLM families and SFT datasets. Our analyses further highlight its versatility and scalability.
Limitation
Due to computational constraints, we do not apply SFTMix to LLM pre-training or instruction-tune models larger than 14B. We also acknowledge that SFTMix requires an initial fine-tuning round to estimate training dynamics-based confidence. While this adds computational cost, it is comparable to prior data selection methods such as LESS (Xia et al., 2024) and Rho-1 (Lin et al., 2024), which also depend on reference models. To mitigate this cost, we explore using confidence scores from weaker LLMs in §5.1, though further experiments are needed to robustly establish their generalization. We also examine SFTMix’s compatibility with lightweight data selection methods in §5.5, but additional evaluation is required to validate its effectiveness across varied strategies.
Appendix A Additional Experiment Results
A.1 Hyperparameter Search
@@
In SFTMix, we use the hyperparameter $\mu$ to control the regularization effect in the training loss $\ell_{\text{SFTMix}}$ and $\alpha$ to control the sampling distribution of $\lambda$. To explore their impacts, we experiment with $\mu\in\{0.05,0.1,0.2,0.5,1\}$ and $\alpha\in\{0.1,0.2,0.5,0.7,1\}$ by instruction-tuning Llama on Alpaca-52K. As shown in Table 9, $\mu=0.2$ achieves the highest performance across most metrics, except for the multi-turn conversational ability in MT-Bench. Similarly, Table 10 shows that $\alpha=0.5$ yields the best overall score on MT-Bench and the highest LC WR on AlpacaEval-2. Hence, we set $\mu=0.2$ and $\alpha=0.5$ in §4 and §5. We also observe that performance remains stable for $\mu$ between $0.1$ and $0.2$ and $\alpha$ between $0.2$ and $0.5$. However, performance declines when $\alpha$ approaches $0.1$ or $1$, where the Beta distribution becomes too flat and the interpolated examples deviate excessively from either endpoint.
In SFTMix, we apply Mixup regularization by interpolating between the most and least confident halves of the dataset. We adopt an equal split for its simplicity and ease of implementation, avoiding additional hyperparameters. To examine this design choice, we also fine-tune Llama-3.1-8B using Mixup between the most and least confident thirds. As shown in Table 11, the equal split continues to yield the largest improvement over NTP. These results suggest that equal partitioning serves as a strong default, while more adaptive strategies remain a promising direction for future work.
A.2 Human Evaluation on AlpacaEval-2
@@
11 $4.7050$ 4.1075 $4.4062$ $3.9450$ $8.2856$
Table 9: Hyperparameter search for $\mu$. We set $\mu=\underline{0.2}$ as the default in SFTMix, as it achieves the highest performance across most metrics.
$\boldsymbol{\alpha}$ MT-Bench AlpacaEval-2
ST MT Overall WR LC WR
$0.1$ $5.1632$ $3.8991$ $4.5312$ $4.8813$ $10.0259$
$0.2$ 5.2237 $3.9103$ $4.5670$ 4.9112 $10.1865$
$\underline{0.5}$ $5.2125$ 3.9525 4.5825 $4.9031$ 10.3195
$0.7$ $5.1209$ $3.9291$ $4.5250$ $4.7912$ $9.8247$
11 $5.0826$ $3.9200$ $4.5013$ $4.7526$ $9.9835$
Table 10: Hyperparameter search for $\alpha$. We set $\mu=\underline{0.5}$ as the default in SFTMix, as it achieves the highest performance across most metrics.
Recipe Mixup MT-Bench AlpacaEval-2
ST MT Overall WR LC WR
NTP - $4.9100$ $3.8150$ $4.3625$ $4.0714$ $8.6528$
SFTMix Half 5.2125 $3.9525$ 4.5825 4.9031 10.3195
SFTMix Third $5.1450$ 3.9850 $4.5650$ $4.2073$ $8.8978$
Table 11: Evaluation of the Mixup ratio in SFTMix. By default, SFTMix applies Mixup between the most and least confident halves, yielding greater improvement.
A.3 Qualitative Example from MT-Bench
Figure 4 presents a test example from the extraction category in MT-Bench, showing responses generated by Llama-3.1-8B instruction-tuned on Alpaca-52K using either NTP or SFTMix. In this example, the LLM trained with SFTMix accurately interprets the instructions from both the first- and second-turn queries, correctly extracting the relevant information from the prompt. Notably, it succeeds in answering the second-turn query, which involves calculating the profit margin and performing a ratio comparison. In contrast, the LLM trained with NTP struggles to differentiate between revenue and profit, leading to incorrect responses in both turns.
Figure 4: A qualitative example from the extraction category in MT-Bench. Compared to its NTP-tuned counterpart, Llama instruction-tuned by SFTMix accurately interprets the queries from both turns and correctly extracts the relevant information from the prompt.
Language Recipe Win Tie Loss
Chinese NTP $30.4\%$ $35.6\%$ $34.0\%$
SFTMix $\textbf{38.8}\%$ $33.2\%$ $28.0\%$
French NTP $34.5\%$ $24.5\%$ $41.0\%$
SFTMix $\textbf{39.0}\%$ $22.5\%$ $38.5\%$
Table 12: Multilingual evaluation of SFTMix. The method consistently improves win rates in both Chinese and French instruction tuning.
A.4 Multilingual Instruction Tuning
To assess SFTMix in multilingual settings, we fine-tune Llama-3.1-8B on the Chinese (Peng et al., 2023) and French (Chen et al., 2023b) variants of Alpaca-52K, following the setup in §4.1. We then compare the resulting models with human-annotated baselines from Aya (Singh et al., 2024) in the respective languages, using GPT-4o (Hurst et al., 2024) as the judge. As shown in Table 12, SFTMix increases the win rate in both languages, demonstrating its cross-lingual generalizability.